# ArtNet: Hierarchical Clustering-Based Artificial Netlist Generator for ML and DTCO Applications

### v1

@@ I-B Related Work
Recently, ML-based DAG generation methods have been proposed [32] [2] [23]. For example, LayerDAG [23] proposes a layerwise autoregressive diffusion model that captures both directional and logical dependencies in DAGs, enabling the conditional generation of synthetic task graphs for hardware benchmarking. By conditioning on hardware metrics (e.g., runtime, resource usage), LayerDAG generates task graphs that can be used to train surrogate models for performance prediction without requiring actual hardware execution. However, since LayerDAG can only generate relatively small graphs (up to 400 nodes), it has limitations in modeling modern VLSI gate-level netlist hypergraphs that have thousands to millions of nodes. LLM4Netlist [31] proposes a step-wise generation framework that translates natural language descriptions into Berkeley logic interchange format (BLIF) netlist s using a fine-tuned LLM. While it achieves state-of-the-art functional correctness on benchmark datasets, it still faces challenges in scaling to highly complex or large-scale circuits, and the inference time grows significantly with circuit size due to iterative decoding. Furthermore, since detailed functional descriptions are required for training, large-scale data generation remains difficult. Such ML-based methods also rely heavily on reference circuits and incur high training costs, making them resource-intensive.
@@ II-C Artificial Netlist Generation Methods
Target design-based. ArtNet also supports generation of artificial netlists from real designs by extracting structural and timing parameters directly from the netlist. This allows the generated netlist to reflect the key characteristics of the original design. As a result, unsharable industrial designs can be abstracted and shared with the academic community, enabling design benchmarking and EDA tool evaluation. When a real netlist is provided, we extract the parameters listed in Table II using OpenDB ($N_{inst}$, $N_{PI}$, $N_{PO}$, $N_{macro}$, $T_{avg}$, $S_{ratio}$), OpenSTA ($D_{min}$, $D_{max}$), and RentCon ($p$, $R_{ratio}$). These parameters are used to generate a SpecFile, which is then used by ArtNet to produce an artificial netlist that statistically resembles the original design (Sections IV-C and IV-E). To make this functionality accessible to a wide range of users, this SpecFile generation flow is implemented in the OpenROAD project [35] as the write_artnet_spec command in the OpenROAD application.
@@ III Our Approach
1
2/* Read input parameters */
3 if Logical hierarchy is present then
4
5 for $i\leftarrow 0;i<\#submodules_{top};i\leftarrow i+1$ do
6
7    $X_{in}^{sub_{i}}\leftarrow$ Read parameters of submodules from SpecFile
8
9
[…]
11Read .lef using OpenDB and map instances to masters
12for each instance or submodule in SpecFile do
13
14 if item is combinational, macro or a submodule then
15
16    $Q_{size}\leftarrow$ Enqueue item
17 else
18
19    $Q_{seq}\leftarrow$ Enqueue item (used in Algorithm 4)
20
21
22/* Hierarchical clustering */
23
[…]
27/* Net generation */
28
[…]
30for $i\leftarrow level_{\text{max}}-1;i>level_{0};i\leftarrow i-1$ do
31
32 $clust_{i}\leftarrow$ cluster corresponding to level $i$
33 if $clust_{i}$ == submodulesubmodule then
34
35     Hierarchical clustering using Algorithm 2
36
37  Generate nets between $clust_{i}.leftChild$ and $clust_{i}.rightChild$ using Algorithm 3
38
39/* PI/PO matching */
40
41if $I_{clust_{root}}\neq N_{PI}||O_{clust_{root}}\neq N_{PO}$ then
42
43  Match $I_{clust_{root}}$ and $O_{clust_{root}}$ to $N_{PI}$ and $N_{PO}$
44
45Netlist $\leftarrow$ Write netlist
@@ III-A Hierarchical Clustering
1
2while $|Q_{size}|\geq 1$ do
3
4 $clust_{A},clust_{B}\leftarrow$ Dequeue from $Q_{size}$
5 $\tau\leftarrow(T(clust_{A})>MaxT(|clust_{B}|))\lor(T(clust_{B})>MaxT(|clust_{A}|))$
6 if $\tau$ then
7
8    $Q_{size}\leftarrow$ Enqueue $clust_{A}$ and $clust_{B}$
9 else
10
11    $clust_{AB}\leftarrow$ Merge $clust_{A}$ and $clust_{B}$
12    $clust_{AB}.rightChild\leftarrow clust_{A}$
13    $clust_{AB}.leftChild\leftarrow clust_{B}$
14    $Q_{size}\leftarrow$ Enqueue $clust_{AB}$
15
[…]
@@ 14    $Q_{size}\leftarrow$ Enqueue $clust_{AB}$
Lines 1-3: We start with $Q_{size}$ that contains initial clusters, and dequeue two clusters $clust_{A}$ and $clust_{B}$. The combining condition ($\tau$) determines whether it is feasible to merge the two clusters, with $T(clust_{A})$ and $T(clust_{B})$ denoting the number of terminals associated with clusters $clust_{A}$ and $clust_{B}$, respectively. The thresholds $MaxT(|clust_{B}|)$ and $MaxT(|clust_{A}|)$ indicate the maximum allowable number of terminals for merging the clusters, and are determined from the Rent’s rule expression:
@@
Lines 4-12: If the combining condition is satisfied, $clust_{A}$ and $clust_{B}$ are merged into a new cluster, $clust_{AB}$. We designate $clust_{A}$ as the right child and $clust_{B}$ as the left child of $clust_{AB}$, capturing the hierarchical relationship between the clusters. Finally, the merged cluster $clust_{AB}$ is enqueued back into $Q_{size}$ for the next iteration. After the clustering is done, the root cluster $clust_{root}$ is dequeued from $Q_{size}$ and returned.
@@ III-B Net Generation
Fig. 4: Net generation flow: (a) calculate In/Out terminals to be created; (b) determine additional FFs from $Q_{seq}$; and (c) make connections.
@@
1
2/* Calculate input and output counts */
3
@@
5/* Calculate number of nets */
6 $clust_{A}\leftarrow clust_{AB}.leftChild$ and $clust_{B}\leftarrow clust_{AB}.rightChild$
7if $(O_{A}+O_{B}-O_{AB})>(I_{A}+I_{B}-I_{AB})$ then
8
9 $|\mathcal{I}_{net}(AB)|\leftarrow(O_{A}+O_{B}-O_{AB}+I_{A}+I_{B}-I_{AB})/2$
10else
11
12 $|\mathcal{I}_{net}(AB)|\leftarrow(O_{A}+O_{B}-O_{AB})$
13
[…]
15/* Make external nets */
16
17Sort($I(A),I(B)$) and Sort($O(A),O(B)$)
18while $k<|\mathcal{I}_{net}(AB)|+|\mathcal{E}_{net}(AB)|$ do
19
20 if Connect $O(A)$ to $I(B)$ then
21
22    $O(A).sinks\leftarrow I(B).sinks$
23     Delete $I(B)$ and kk++
24 else
25
26    $I(B)\rightarrow I(B)_{i+1}$
27
28 if Connect $O(B)$ to $I(A)$ then
29
30    $O(B).sinks\leftarrow I(A).sinks$
31     Delete $I(A)$ and kk++
32 else
33
34    $I(A)\rightarrow I(A)_{i+1}$
35
36 if no connection available then
37
38     break
39
40
41/* Make internal nets */
42
[…]
44while $|\mathcal{I}_{net}(AB)|>0$ do
45
46 $\mathcal{I}_{net}(AB)\leftarrow\mathcal{E}_{net}(AB)$
47 $|\mathcal{I}_{net}(AB)|=|\mathcal{I}_{net}(AB)|-1$
48
49Delete $clust_{A}$ and $clust_{B}$
@@ 46 $\mathcal{I}_{net}(AB)\leftarrow\mathcal{E}_{net}(AB)$
Then, we get the number of input terminals $I_{C}$ by T(1G)T*(1-G) and the number of output terminals $O_{C}$ by TGT*G.
Lines 3-9: Let clusters $A$ and $B$ have sets of input/output nets $I_{A}$$O_{A}$ and $I_{B}$$O_{B}$, respectively. If the number of output nets exceeds the number of input nets, then the internal net count ($\mathcal{I}_{net}$) is set to the average of these values, balancing connections between inputs and outputs. Otherwise, it is set to the output net count. The external net count is then calculated by subtracting the input net count from the total number of input nets, ensuring that the sum of internal and external net counts matches the module’s connection capacity. This approach maintains a balanced connection distribution.
Lines 10-24: We sort output nets by the logic depth of their sink instances (from largest to smallest), and input nets by the logic depth of their source instances (from smallest to largest). When connecting these nets, we ensure that no resulting path depth exceeds $D_{max}$ (if a combined path length exceeds $D_{max}$, the connection is skipped). If cluster $A$’s output net $O(A)$ can connect to cluster $B$’s input net $I(B)$, we remove $I(B)$ and connect it to $O(A)$. Conversely, if $O(B)$ can connect to $I(A)$, we remove $I(A)$ and connect it to $O(B)$. If neither of these connections is possible, we exit the loop (Line 19). This approach prioritizes the longest paths first, addressing the most critical paths early in the process.
Lines 25-31: External nets are sorted in descending order of depth, so as to prevent large-depth paths from persisting through subsequent clustering stages. After sorting, they are converted to internal nets until the required number of internal nets is reached. Last, we remove $clust_{A}$ and $clust_{B}$ and return the merged cluster $clust_{C}$.
@@
$FF_{new}=\lfloor S_{ratio}\times|{clust}_{C}|-(FF_{{clust}_{A}}+FF_{{clust}_{B}})\rfloor.$ (4)
@@
1
2/* Prepare a flip-flop instance */
3
4if $Q_{seq}$ is not empty then
5 $FF\leftarrow$ Dequeue from $Q_{seq}$
6else
7 $FF\leftarrow$ default flip-flop FFFF
8
9Update $S_{ratio}$, $N_{inst}$ of $clust_{AB}$
10/* Insert $F$ between the net’s source and sinks */
11
12Create a net $Net_{D}$ to drive the data input $D$ of FFFF
13Move the source of $O(A)$ to $Net_{D}$
14Connect the data output $Q$ of FFFF to the source of $O(A)$
15Add $Net_{D}$ to $\mathcal{I}_{net}(A)$
@@ 7 $FF\leftarrow$ default flip-flop FFFF
Lines 1-5: We prepare a flip-flop instance $FF$ by either dequeuing from the sequential queue $Q_{seq}$ or instantiating a default flip-flop when the queue is empty. The default flop is determined either by a user-specified flip-flop or by selecting the most frequently used flip-flop in the design.
@@ III-C PI/PO Matching
Fig. 5: PI/PO matching methods: (a) add PI; (b) delete PI; (c) add PO; and (d) delete PO.
@@ IV-A Runtime Efficiency
To obtain the data in the figure, we sweep instance counts from 500,000 to 100M with step size 500,000, and sweep $p$ from 0.45 to 0.55 with step size 0.05. Other parameters are fixed: $S_{ratio}$ is set to 0.2, $G_{avg}$ is set to 0.3, and minimum and maximum logic depths are fixed at 1 and 40, respectively, as their impact on complexity is negligible. ArtNet requires between 1,731 and 3,491 seconds of runtime to generate 100M-instance netlists with these parameter settings.
Fig. 6: Runtime analysis: (a) Runtime vs. #Insts; and (b) Runtime breakdown.

### v2

@@ I-B Related Work
Recently, ML-based DAG generation methods have been proposed [34] [2] [25]. For example, LayerDAG [25] proposes a layerwise autoregressive diffusion model that captures both directional and logical dependencies in DAGs, enabling the conditional generation of synthetic task graphs for hardware benchmarking. By conditioning on hardware metrics (e.g., runtime, resource usage), LayerDAG generates task graphs that can be used to train surrogate models for performance prediction without requiring actual hardware execution. However, since LayerDAG can only generate relatively small graphs (up to 400 nodes), it has limitations in modeling modern VLSI gate-level netlist hypergraphs that have thousands to millions of nodes. LLM4Netlist [33] proposes a step-wise generation framework that translates natural language descriptions into Berkeley logic interchange format (BLIF) netlists using a fine-tuned LLM. While it achieves state-of-the-art functional correctness on benchmark datasets, it still faces challenges in scaling to highly complex or large-scale circuits, and the inference time grows significantly with circuit size due to iterative decoding. Furthermore, since detailed functional descriptions are required for training, large-scale data generation remains difficult. HYGENE [9] enables the generation of hyperedges using a hypergraph generation method. However, it also has limitations in generating large graphs with multi-million nodes. Overall, ML-based methods tend to rely heavily on reference circuits and incur high training costs, making them resource-intensive ; moreover, they have super-linear time complexity with the number of nodes [10].
@@ II-C Artificial Netlist Generation Methods
Target design-based. ArtNet also supports generation of artificial netlists from real designs by extracting structural and timing parameters directly from the netlist. This allows the generated netlist to reflect the key characteristics of the original design. As a result, unsharable industrial designs can be abstracted and shared with the academic community, enabling design benchmarking and EDA tool evaluation. When a real netlist is provided, we extract the parameters listed in Table II using OpenDB ($N_{inst}$, $N_{PI}$, $N_{PO}$, $N_{macro}$, $T_{avg}$, $S_{ratio}$), OpenSTA ($D_{min}$, $D_{max}$), and RentCon ($p$, $R_{ratio}$). These parameters are used to generate a SpecFile, which is then used by ArtNet to produce an artificial netlist that statistically resembles the original design (Sections IV-C and IV-E). To make this functionality accessible to a wide range of users, this SpecFile generation flow is implemented in the OpenROAD project [37] as the write_artnet_spec command in the OpenROAD application , and a more detailed SpecFile description is provided in [42].
@@ III Our Approach
/* Read input parameters */
if Logical hierarchy is present then
 for $i\leftarrow 0;i<\#submodules_{top};i\leftarrow i+1$ do
    $X_{in}^{sub_{i}}\leftarrow$ Read parameters of submodules from SpecFile
   end for
end if
$X_{in}^{top}\leftarrow$ Read parameters of top module from SpecFile
Read .lef using OpenDB and map instances to masters
for each instance or submodule in SpecFile do
 if item is combinational, macro or a submodule then
    $Q_{size}\leftarrow$ Enqueue item
   end if
 else
    $Q_{seq}\leftarrow$ Enqueue item (used in Algorithm 4)
   end if
end for
/* Hierarchical clustering */
$\tau\leftarrow$ Generate combining constraints from $X_{in}^{top}$, as defined in Algorithm 2
$clust_{root}\leftarrow$ Get root cluster from clustering using Algorithm 2
$H(V,E)\leftarrow$ Generate hierarchy tree from $clust_{root}$
/* Net generation */
$level_{\text{max}}\leftarrow$ $H(V,E)$
for $i\leftarrow level_{\text{max}}-1;i>level_{0};i\leftarrow i-1$ do
 $clust_{i}\leftarrow$ cluster corresponding to level $i$
 if $clust_{i}$ == submodulesubmodule then
     Hierarchical clustering using Algorithm 2
   end if
  Generate nets between $clust_{i}.leftChild$ and $clust_{i}.rightChild$ using Algorithm 3
end for
/* PI/PO matching */
if $I_{clust_{root}}\neq N_{PI}||O_{clust_{root}}\neq N_{PO}$ then
  Match $I_{clust_{root}}$ and $O_{clust_{root}}$ to $N_{PI}$ and $N_{PO}$
end if
Netlist $\leftarrow$ Write netlist
@@ III-A Hierarchical Clustering
$fail\_count\leftarrow 0$
$fail\_limit\leftarrow|Q_{size}|$
while $|Q_{size}|\geq 1$ do
 $clust_{A},clust_{B}\leftarrow$ Dequeue from $Q_{size}$
 $\tau\leftarrow(T(clust_{A})>MaxT(|clust_{B}|))\lor(T(clust_{B})>MaxT(|clust_{A}|))$
 if $\tau$ then
    $Q_{size}\leftarrow$ Enqueue $clust_{A}$ and $clust_{B}$
     $fail\_count\leftarrow fail\_count+1$
    if $fail\_count\geq fail\_limit$ then
        $\sigma_{T}\leftarrow\sigma_{T}*1.2$
        $fail\_count\leftarrow 0$
        $fail\_limit\leftarrow|Q_{size}|$
      end if
   end if
 else
    $clust_{AB}\leftarrow$ Merge $clust_{A}$ and $clust_{B}$
    $clust_{AB}.rightChild\leftarrow clust_{A}$
    $clust_{AB}.leftChild\leftarrow clust_{B}$
    $Q_{size}\leftarrow$ Enqueue $clust_{AB}$
     $fail\_count\leftarrow 0$
     $fail\_limit\leftarrow|Q_{size}|$
   end if
end while
$clust_{root}\leftarrow$ Dequeue from $Q_{size}$
@@
Lines 1-12: We start with $Q_{size}$ that contains initial clusters, and initialize a failure counter along with a threshold based on the queue size. We dequeue two clusters $clust_{A}$ and $clust_{B}$. The combining condition ($\tau$) determines whether it is feasible to merge the two clusters, with $T(clust_{A})$ and $T(clust_{B})$ denoting the number of terminals associated with clusters $clust_{A}$ and $clust_{B}$, respectively. If $\tau$ holds, we re-enqueue $clust_{A}$ and $clust_{B}$ and increment fail_countfail\_count; when it reaches fail_limitfail\_limit, we relax the merge criterion by updating $\sigma_{T}$. The thresholds $MaxT(|clust_{B}|)$ and $MaxT(|clust_{A}|)$ indicate the maximum allowable number of terminals for merging the clusters, and are determined from the Rent’s rule expression:
@@
Lines 13-21: If the combining condition is satisfied, $clust_{A}$ and $clust_{B}$ are merged into a new cluster, $clust_{AB}$. We designate $clust_{A}$ as the right child and $clust_{B}$ as the left child of $clust_{AB}$, capturing the hierarchical relationship between the clusters. Finally, the merged cluster $clust_{AB}$ is enqueued back into $Q_{size}$ for the next iteration. After the clustering is done, the root cluster $clust_{root}$ is dequeued from $Q_{size}$ and returned.
@@ III-B Net Generation
Fig. 4: Net generation flow: (a) calculate In/Out terminals to be created; (b) determine additional FFs from $Q_{seq}$; and (c) make connections. (Refer to Table III for notation.)
@@
/* Calculate input and output counts */
$I_{AB},O_{AB}\leftarrow$ $getIO(|clust_{AB}|)$
/* Calculate number of nets */
$clust_{A}\leftarrow clust_{AB}.leftChild$ and $clust_{B}\leftarrow clust_{AB}.rightChild$
if $(O_{A}+O_{B}-O_{AB})>(I_{A}+I_{B}-I_{AB})$ then
 $|\mathcal{I}_{net}(AB)|\leftarrow(O_{A}+O_{B}-O_{AB}+I_{A}+I_{B}-I_{AB})/2$
end if
else
 $|\mathcal{I}_{net}(AB)|\leftarrow(O_{A}+O_{B}-O_{AB})$
end if
$|\mathcal{E}_{net}(AB)|\leftarrow I_{A}+I_{B}-I_{AB}-|\mathcal{I}_{net}(AB)|$
Sort($I(A),I(B)$) and Sort($O(A),O(B)$)
Initialize indices $i\leftarrow 0$, $j\leftarrow 0$, and counter $k\leftarrow 0$
/* Make external (inter-cluster) nets */
while $k<|\mathcal{I}_{net}(AB)|+|\mathcal{E}_{net}(AB)|$ do
  $Cond_{A\to B}\leftarrow(Depth(O(A)_{0})+Depth(I(B)_{i})\leq D_{max})$
  $Cond_{B\to A}\leftarrow(Depth(O(B)_{0})+Depth(I(A)_{j})\leq D_{max})$
 if $Cond_{A\to B}$ then
     $O(A)_{0}\leftarrow Connect(O(A)_{0},I(B)_{i})$
     Delete $I(B)_{i}$, Append $O(A)_{0}$ to $\mathcal{E}_{net}(AB)$, and kk++
   end if
 else
     $i\leftarrow i+1$
   end if
 if $Cond_{B\to A}$ then
     $O(B)_{0}\leftarrow Connect(O(B)_{0},I(A)_{j})$
     Delete $I(A)_{j}$, Append $O(B)_{0}$ to $\mathcal{E}_{net}(AB)$, and kk++
   end if
 else
     $j\leftarrow j+1$
   end if
 if no connection available then
     break
   end if
end while
/* Convert external nets to internal nets */
@@
while $|\mathcal{I}_{net}(AB)|>0$ do
  Append $\mathcal{E}_{net}(AB)_{0}$ to $\mathcal{I}_{net}(AB)$
  Remove $\mathcal{E}_{net}(AB)_{0}$ from $\mathcal{E}_{net}(AB)$ and $|\mathcal{I}_{net}(AB)|$--
end while
Delete $clust_{A}$ and $clust_{B}$
@@
Then, we get the number of input terminals $I_{AB}$ by T(1G)T*(1-G) and the number of output terminals $O_{AB}$ by TGT*G.
Lines 3-9: Let clusters $A$ and $B$ have sets of input/output nets $I_{A}$$O_{A}$ and $I_{B}$$O_{B}$, respectively. If the number of output nets exceeds the number of input nets, then the internal net count ($\mathcal{I}_{net}$) is set to the average of these values, balancing connections between inputs and outputs. Otherwise, it is set to the output net count. The external net count is then calculated by subtracting the input net count from the total number of input nets, ensuring that the sum of internal and external net counts matches the module’s connection capacity. This approach maintains a balanced connection distribution , and multi-sink nets naturally arise when terminal counts are matched, since a parent-level cluster’s terminal aggregates multiple underlying sinks.
Lines 10-27: We sort output nets by logic depth , the logic cell hop count to the farthest adjacent sequential cell, in descending order, and input nets by logic depth in ascending order. $I(A)$ and $O(A)$ denote the input/output terminal lists of cluster $A$. $I(A)_{i}$ denotes the $i^{th}$ element of $I(A)$ and $I(A)_{0}$ refers to the first element of $I(A)$. We maintain indices $i$ and $j$ as pointers into $I(B)$ and $I(A)$, respectively, and use $k$ as a counter for the number of inter-cluster connections. We initialize $i,j,k$ to 0. When connecting these nets, we ensure that no resulting path depth exceeds $D_{max}$ (if a combined path length exceeds $D_{max}$, the connection is skipped). We check whether connecting $O(A)_{0}$ to $I(B)_{i}$ or $O(B)_{0}$ to $I(A)_{j}$ satisfies the depth constraint. When it is satisfied, the corresponding input net is removed and the resulting output net is appended to the external net set $\mathcal{E}_{net}$. If neither of these connections is possible, we exit the loop ( Line 27). This approach prioritizes the longest paths first, addressing the most critical paths early in the process.
Lines 28-34: External nets are sorted in descending order of their corresponding input nets’ depth to prevent large-depth paths from persisting through subsequent clustering stages. After sorting, they are converted to internal nets until the required number of internal nets is reached. Last, we remove $clust_{A}$ and $clust_{B}$ and return the merged cluster $clust_{AB}$.
@@
[…]
@@
/* Prepare a flip-flop instance */
if $Q_{seq}$ is not empty then
 $FF\leftarrow$ Dequeue from $Q_{seq}$
else
 $FF\leftarrow$ default flip-flop FFFF
end if
Update $S_{ratio}$, $N_{inst}$ of $clust_{AB}$
/* Insert $F$ between the net’s source and sinks */
Create a net $Net_{D}$ to drive the data input $D$ of FFFF
Move the source of $O(A)$ to $Net_{D}$
Connect the data output $Q$ of FFFF to the source of $O(A)$
Add $Net_{D}$ to $\mathcal{I}_{net}(A)$
@@
Lines 1-5: We prepare a flip-flop instance $FF$ by either dequeuing from the sequential queue $Q_{seq}$ or instantiating a default flip-flop when the queue is empty. The default flip-flop is determined either by a user-specified flip-flop or by selecting the most frequently used flip-flop in the design.
@@ III-C PI/PO Matching
Fig. 5: PI/PO matching methods.
@@ IV-A Runtime Efficiency
To obtain the data in the figure, we sweep instance counts from 500,000 to 100M with step size 500,000, and sweep $p$ from 0.45 to 0.55 with step size 0.05. Other parameters are fixed: $S_{ratio}$ is set to 0.2, $G_{avg}$ is set to 0.3, and minimum and maximum logic depths are fixed at 1 and 40, respectively, as their impact on complexity is negligible. We cap the maximum runtime at 3 hours; under this constraint, ANG generates netlists up to 250K instances, while GNL succeeds up to 50.5M instances. ArtNet requires between 980 and 3,212 seconds of runtime to generate 100M-instance netlists with these parameter settings , while GNL requires a minimum of 29,000 seconds of runtime to generate 100M-instance netlists and ANG requires a minimum of 38,000 seconds of runtime to generate 500K-instance netlists. This runtime gap is primarily due to ANG’s flat, distribution-matching–based graph construction: edges are inserted one by one to reduce mismatches to target statistics (e.g., fan-in/fan-out distributions), and each insertion requires a global search over candidate node pairs. As design size increases, the candidate space grows rapidly, resulting in strongly superlinear runtime scaling.
Fig. 6: Runtime analysis: (a) Runtime vs. #Insts; and (b) ArtNet Runtime breakdown.